Prove: 1-(sec^6a+cos^6a)=3sin^2acos^2a
1
Expert's answer
2011-10-20T08:46:57-0400
To prove the equality means to show that it holds for every value of a. In this case it's not true
For example. we can substitute a=pi/4:
left part& 1-((sec pi/4)^6+(cos pi/4)^6)=1-((sqrt(2))^6+1/(sqrt(2))^6)=1-(8+1/8)=-7-1/8=-57/8
right part 3*(sin pi/4)^2 * (cos pi/4)^2=3*1/2*1/2=3/4
-57/8 = 3/4 that's false so given equality doesn't hold for a=pi/4 . It means that equality is not true for every a